OpenStudy (anonymous):

Find all the complex roots. Write the answer in the indicated form. The complex cube roots of 27(cos 234° + i sin 234°) (polar form) B. 3(cos 78° + i sin 78°), 3(cos 118° + i sin 118°), 3(cos 158° + i sin 158°) C. -3(cos 78° + i sin 78°), 3(cos 118° + i sin 118°), -3(cos 158° + i sin 158°) D. 3(cos 78° + i sin 78°), 3(cos198° + i sin 198°), 3(cos 318° + i sin 318°)

OpenStudy (anonymous):

@bibby

zepdrix (zepdrix):

\[\Large\rm \left[27\left(\cos234^o+\mathcal i \sin234^o\right)\right]^{1/3}\]So we need to apply our De Moivre's Theorem, yes? :)

zepdrix (zepdrix):

We should think of our angle like this:\[\Large\rm \left[27\left(\cos\left(234^o+360^ok\right)+\mathcal i \sin\left(234^o+360^ok\right)\right)\right]^{1/3}\]We can get that same angle by spinning around.

zepdrix (zepdrix):

And our roots will correspond to k=0,1,2. The first three k values. That will give us our three unique roots.

zepdrix (zepdrix):

\[\Large\rm 3\left[\cos\left(234^o+360^ok\right)+\mathcal i \sin\left(234^o+360^ok\right)\right]^{1/3}\]Take the cube root of the 27.

zepdrix (zepdrix):

Remember how to apply your De Moivre's Theorem? We just multiply the angles by this exponent.\[\Large\rm =3\left[\cos\left(\frac{234^o+360^ok}{3}\right)+\mathcal i \sin\left(\frac{234^o+360^ok}{3}\right)\right]\]So we're multiplying by 1/3, or dividing by 3.

zepdrix (zepdrix):

Plug in k=0 to get your first solution. k=1 for the next, and k=2 for the final solution.

zepdrix (zepdrix):

Lady Skitttleeeee :U where you at got it?

OpenStudy (anonymous):

Where do I plug in k=0? I don't understand any of this, it really confuses me

zepdrix (zepdrix):

So you're at some angle 234 degrees.

zepdrix (zepdrix):

But we can also get to that same location by spinning around the circle, yah?